OpenStudy (anonymous):

write the equation of a line in the form ax + by = c that passes through (-5, 3); m= 3/4

OpenStudy (anonymous):

y = (3/4)x + b 3 = (3/4)(-5) + b b = 27/4 y = (3/4)x + (27/4) -(3/4)x +1y = (27/4)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem